# Gaussian-Det: Learning Closed-Surface Gaussians for 3D Object Detection

### v1

@@ 4.2 Main Results
Table 1: Quantitative results on the 3D-FRONT and ScanNet datasets. $\dagger$: we report the results of NeRF-MAE which involves the pre-training on 3D-FRONT and two other exterior databases, including HM3D Ramakrishnan et al. [2021] and Hypersim Roberts et al. [2021].
@@
G-VoteNet Qi et al. [2019] 81.5 61.6 73.0 49.6 78.5 34.2 66.8 18.2
G-GroupFree Liu et al. [2021] 84.9 63.7 72.1 45.1 75.2 37.6 60.1 20.4
G-FCAF3D Rukhovich et al. [2022a] 89.1 56.9 73.1 35.2 90.2 42.4 63.7 18.5
G-BRNet Xu et al. [2022b] 89.7 75.3 88.2 71.0 71.1 32.2 63.1 19.3
@@
Compared Methods: We have summarized the quantitative results on 3D-FRONT and ScanNet in Table 1. For the comparison with the pure multi-view based methods, we follow Irshad et al. [2024] to include ImVoxelNet [Rukhovich et al., 2022b], NeRF-RPN [Hu et al., 2023] and NeRF-MAE [Irshad et al., 2024] in Table 1. ImVoxelNet directly extracts 3D volumetric features projected from 2D image planes. The NeRF-based methods [Hu et al., 2023, Irshad et al., 2024] apply 3D convolutions on the uniformly sampled points in the radiance fields. Note that NeRF-MAE includes the pre-training on extra datasets. Moreover, to demonstrate the effectiveness of Gaussian-Det more intuitively, we experimented on the point cloud based detectors [Qi et al., 2019, Liu et al., 2021, Rukhovich et al., 2022a, Cheng et al., 2021] (prefixed by “G-”) which merely contains the backbone and detection heads in Section 3.2. The upper part of Table 1 are cited from Irshad et al. [2024], for which we have provided a more in-depth comparison in the Appendix pages.
@@
Failure Case: As for illustrative comparison, we firstly present the failure case by Gaussian-Det as shown in Figure 5. Despite focusing on more informative surface representations, Gaussian-Det may still struggle at handling detecting the object with severe occlusion, which is an inherent challenge for current multi-view based 3D object detectors. Moreover, since the surfaces of the objectness that contains jointed objects can also be regarded as closed, Gaussian-Det may mistake them as a whole. Please refer to the Appendix pages for more failure cases.
Qualitative Results: We then display the qualitative comparisons in Figure 6, where we set the 2D rendered results of the input Gaussians as visual reference. We can see that compared with ground truth bounding boxes, Gaussian-Det is capable of preventing redundant object predictions that are potentially caused by outliers, leading to higher detection accuracy. Figure 7 provides a more in-depth visualization on the discriminative objectness deduction of Gaussian-Det. For each detected bounding box, we have marked the corresponding flux value computed in Eqn.(10) that measures the surface closure. Predictions whose absolute flux values are larger or less than one are marked with blue or red respectively. We can see that in the presented examples, a favourable detection well corresponds to a relatively higher degree of closure (lower absolute value of flux). With the holistic surface coalescence by CIM, Gaussian-Det is aware of the reliability of proposals and able to refine them for higher accuracy. Please refer to the Appendix pages for more qualitative results.
Figure 6: Qualitative Results on 3D-FRONT (top two rows) and ScanNet (bottom two rows).
@@ 4.3 Ablation Studies and Analysis
Please refer to the Appendix pages for additional ablational experiments, including the choices of hyper-parameters and the comparison on learning curves before and after incorporating the probabilistic feature residual.
@@
Comparison with Point-based Partial Surface Modeling: It is worth mentioning that RepSurf [Ran et al., 2022], a point cloud based convolutional operator, also considers the surface structure within local regions. We accordingly replaced the backbone of G-BRNet [Cheng et al., 2021] with the umbrella variant of RepSurf. The results on ScanNet are shown in Table 4. The observation is two-fold. Firstly, the incorporation of RepSurf does augment the point features with informative depiction on local surfaces (+0.4% on AP50). However, RepSurf is prone to be severely distracted by the outlier Gaussians since it only considers the very local surfaces, which harms the detection quality drastically (-15.34% on AP25). Secondly, merely focusing on the very local (umbrella-shaped in this case) surface while ignoring the holistic surface modeling leads to sub-optimal results (-4.8% on AP50). This indicates that the comprehensive surface modeling strategy enables Gaussian-Det to better capture the intricate geometry of objects, thereby contributing to a sizably improved 3D detection accuracy. Such awareness of holistic surface representation is essential to prevent the distraction by outliers.
@@
Description AP25 AP50
G-BRNet 63.1 19.3
+ RepSurf-U 47.7 19.7
@@ 5 Conclusion
Future Works: Since the degree of surface closure of the objectness that contains joint objects is also relatively high, such scenarios might be challenging for Gaussian-Det. Future work can incorporate the color information to discriminate the objects and calculate the flux values separately. In addition, the overall closed surfaces can potentially be an informative clue to rule out the outlier 3D Gaussians, thereby in turn improving the quality of 3D reconstruction.
Appendix A More Technical Details
A.1 Details on Calculating Surface Normal
@@
We define eigenvector with smallest eigenvalue as the Surface Normal $\mathbf{n}_{p}$. The defination is consistent with [Guédon and Lepetit, 2024]. However, $\mathbf{n}_{p}$ is unoriented, meaning it can point either inside or outside. Therefore, we determine the surface normal direction by considering the object center (in practice, we use proposal center). To elaborate, the orientation of surface normal will align with the orientation pointing from the object center to the Gaussian center. An alternative strategy is to follow [Ran et al., 2022] to determine the orientation. Specifically, we ensure that the first element of the normal vector is non-negative. Then we determine the orientation using an instance-level random inversion with $p=50\%$. Ablation study in Table 2 shows that the result of strategy of Ran et al. [2022] is lower than the aforementioned strategy.
A.2 More Details on Training
@@
Appendix B More Ablation Experiments and Analysis
B.1 Hyper-parameter Setup
@@
B.2 Impacts of $F_{V}$ on the Training Process
@@
Appendix C More Visulization Results
C.1 More Qualitative Results under Challenging Scenarios
@@
C.2 More Failure Cases
@@
Appendix D More Detailed Quantitative Comparison
D.1 Detailed Comparison with NeRF-RPN and NeRF-MAE

### v2

@@ 4.2 Main Results
Table 1: Quantitative results on the 3D-FRONT and ScanNet datasets. $\dagger$: we report the results of NeRF-MAE which involves the pre-training on 3D-FRONT and two other exterior databases, including HM3D [Ramakrishnan et al., 2021] and Hypersim [Roberts et al., 2021].
@@
G-VoteNet [Qi et al., 2019] 81.5 61.6 73.0 49.6 78.5 34.2 66.8 18.2
G-GroupFree [Liu et al., 2021] 84.9 63.7 72.1 45.1 75.2 37.6 60.1 20.4
G-FCAF3D [Rukhovich et al., 2022a] 89.1 56.9 73.1 35.2 90.2 42.4 63.7 18.5
G-BRNet [Xu et al., 2022b] 89.7 75.3 88.2 71.0 71.1 32.2 63.1 19.3
@@
Figure 4: Qualitative Results on 3D-FRONT (top two rows) and ScanNet (bottom two rows). We visualize each bounding box with a unique color for clear illustration.
Compared Methods: We have summarized the quantitative results on 3D-FRONT and ScanNet in Table 1. For the comparison with the pure multi-view based methods, we follow [Irshad et al., 2024] to include ImVoxelNet [Rukhovich et al., 2022b], NeRF-RPN [Hu et al., 2023] and NeRF-MAE [Irshad et al., 2024] in Table 1. ImVoxelNet directly extracts 3D volumetric features projected from 2D image planes. The NeRF-based methods [Hu et al., 2023, Irshad et al., 2024] apply 3D convolutions on the uniformly sampled points in the radiance fields. Note that NeRF-MAE includes the pre-training on extra datasets. Moreover, to demonstrate the effectiveness of Gaussian-Det more intuitively, we experimented on the point cloud based detectors [Qi et al., 2019, Liu et al., 2021, Rukhovich et al., 2022a, Cheng et al., 2021] (prefixed by “G-”) which merely contains the backbone and detection heads in Section 3.2. The upper part of Table 1 are cited from [Irshad et al., 2024], for which we have provided a more in-depth comparison in the Appendix pages.
@@
Qualitative Results: As for illustrative comparison, we firstly display the qualitative comparisons in Figure 4, where we set the 2D rendered results of the input Gaussians as visual reference. We can see that compared with ground truth bounding boxes, Gaussian-Det is capable of preventing redundant object predictions that are potentially caused by outliers, leading to higher detection accuracy. Figure 6 provides a more in-depth visualization on the discriminative objectness deduction of Gaussian-Det. For each detected bounding box, we have marked the corresponding flux value computed in Eqn. (9) that measures the surface closure. Predictions whose absolute flux values are larger or less than one are marked with blue or red respectively. We can see that in the presented examples, a favourable detection well corresponds to a relatively higher degree of closure (lower absolute value of flux). With the holistic surface coalescence by CIM, Gaussian-Det is aware of the reliability of proposals and able to refine them for higher accuracy. Please refer to the Appendix pages for more qualitative results under challenging scenarios.
@@
Failure Case: We then present the failure case by Gaussian-Det in Figure 5. Despite focusing on more informative surface representations, Gaussian-Det may still struggle at handling detecting the object with severe occlusion, which is an inherent challenge for current multi-view based 3D object detectors. Moreover, since the surfaces of the objectness that contains jointed objects can also be regarded as closed, Gaussian-Det may mistake them as a whole. Please refer to the Appendix for more failure cases.
@@ 4.3 Ablation Studies and Analysis
Please refer to the Appendix pages for additional ablational experiments, including the choices of hyper-parameters, Gaussian Splatting representation and the comparison on learning curves before and after incorporating the probabilistic feature residual.
@@
Detecting with Noised Poses: To furtherly validate on inferring the objectness from low-quality data, we experimented on adding normally distributed perturbations with the mean of 0 and variance of 0.07 to camera extrinsics when reconstructing them. The corresponding quantitative results are shown in Table 6. We can see that with noised poses, Gaussian-Det showcases less precision decline on both AP and AR, which verifies the effectiveness of Gaussian-Det against either inherent outlier 3D Gaussians or low-quality input poses. Please refer to the Appendix for qualitative comparisons.
Efficacy on 3D Instance Segmentation: We have further validated the generality of our proposed method on the task of open-vocabulary 3D instance segmentation. On the Gaussian-Grouping baseline [Ye et al., 2024], we incorporated the estimation of surface closure as a prior on the reliability of the object instance, thus supporting or suppressing the segmented results. From the results in Table 6, we can see that leveraging the surface closure provides informative prior on objectness deduction and largely enhances the performances of instance segmentation in 3D scenes. The rendering results of qualitative examples containing multiple object instances are shown in Figure 7. It can be seen that the incorporation of the prior of surface closure contributes to fewer noisy predictions off the table. Moreover, it substantially speeds up the training period and saves the GPU memory footprint (see Appendix) compared with the baseline method.
Comparison with Point-based Partial Surface Modeling: We note that RepSurf [Ran et al., 2022], a point cloud based convolutional operator, also considers the surface structure within local regions. We accordingly replaced the backbone of G-BRNet [Cheng et al., 2021] with the umbrella variant of RepSurf. The results on ScanNet are shown in Table 6. RepSurf is prone to be severely distracted by the outlier Gaussians since it only considers the very local surfaces, which harms the detection quality drastically (-15.34% on AP25). Secondly, merely focusing on the very local (umbrella-shaped in this case) surface while ignoring the holistic surface modeling leads to sub-optimal results (-4.8% on AP50). This indicates that the comprehensive surface modeling strategy enables Gaussian-Det to better capture the intricate geometry of objects, thereby contributing to a sizably improved 3D detection accuracy. Such awareness of holistic surface representation is essential to prevent the distraction by outliers.
@@
Method AP25 AP50
G-BRNet [Xu et al., 2022b] 63.1 19.3
+ RepSurf-U [Ran et al., 2022] 47.7 19.7
@@
Table 5: Comparison under the setting of adding noises to camera poses.
Methods Noised Poses AR25 AP25
G-BRNet [Cheng et al., 2021] 89.7 88.2
G-BRNet [Cheng et al., 2021] 74.3 (-15.4) 67.6 (-20.6)
Gaussian-Det (Ours) 97.9 96.7
Gaussian-Det (Ours) 86.8 (-11.1) 81.5 (-15.2)
Table 6: Comparison on 3D instance segmentation.
Methods mIoU mBIoU
DEVA [Cheng et al., 2023] 46.2 45.1
LERF [Kerr et al., 2023] 33.5 30.6
SA3D [Cen et al., 2023] 24.9 23.8
LangSplat [Qin et al., 2024] 52.8 50.5
Gaussian-Grouping [Ye et al., 2024] 69.7 67.9
Gaussian-Grouping + Ours 76.5 (+6.8) 73.3 (+5.4)
Figure 7: Rendered qualitative results of open-vocabulary 3D instance segmentation.
@@ 5 Conclusion
Future Works: Since the degree of surface closure of the objectness that contains joint objects is also relatively high, such scenarios might be challenging for Gaussian-Det. Future work can incorporate the color information to discriminate the objects and calculate the flux values separately. In addition, the exploration of large-scale datasets on autonomous driving is a promising practice.
Acknowledgment: This work was supported in part by the National Natural Science Foundation of China under Grant 62206147.
Appendix A Additional Details on the Datasets
A.1 Dataset Overview
In Figure 8, we visualize the scenes in the 3D-FRONT dataset with ground-truth bounding boxes. Each column shows two different views of the same scene. A total of 159 usable rooms are manually selected, cleaned, and rendered to ensure have high quality for scene reconstruction and object detection, as done by NeRF-RPN [Hu et al., 2023].
Figure 8: Snapshots of scenes in 3D-FRONT and the corresponding 3D bounding box annotations.
A.2 Statistical Details
Figure 9: Statistics on surface closure of two datasets [Fu et al., 2021, Dai et al., 2017]. Most objects in both datasets contain a higher degree of surface closure. The statistics are collected from our reconstructed 3D Gaussians via 3DGS+SuGAR.
Appendix B Additional Technical Details
B.1 More Details on Network Architecture
For corresponding feature $F^{cand}$ generation, we use the same PointNet++ backbone in VoteNet [Qi et al., 2019], which is the backbone $\mathcal{B}$ in Eqn. 4. Specifically, it has four set abstraction (SA) layers and two feature propagation/upsampling (FP) layers. SA layers sub-sample the input to 2048, 1024, 512 and 256 Gaussians respectively. The two FP layers up-sample the 4th SA layer’s output back to 1024 Gaussians with 256-dim features and 3D coordinates.
The querying-and-pooling function in Eqn. 5 and Eqn. 10 has the output size of 128. Therefore, $f^{part}$ and $f^{hol}$ have the same feature channels of 128 and $f^{final}$ has the channel of 256.
The stacked MLP for proposal refinement in Eqn. 12 has output size of 128, 128, 9. In these 9 channels, the first two are for objectness classification, the following one is for heading angle refinement and the last six are for offsets refinement. The initial proposal has the same shape with refined proposal and is generated in the same way.
For the point cloud based methods, the original 3D Gaussians $\mathbb{G}=\{\mathbf{g}_{p}=(\mathbf{\mu}_{p},\mathbf{\Sigma}_{p},\alpha_{p},\mathbf{c}_{p})\}^{P}_{p}$ (generated by 3DGS+SuGAR) were used as the input, where the Gaussian center $\mathbf{\mu}_{p}$ was regarded as the input point-wise coordinate. $\mathbf{\Sigma}_{p},\alpha_{p},\mathbf{c}_{p}$ were concatenated and regarded as the input point-wise features. Except for the first layer whose input dimension was modified to fit the input Gaussians, we faithfully followed the network and training configurations of the point cloud based methods [Qi et al., 2019, Liu et al., 2021, Rukhovich et al., 2022a, Xu et al., 2022b] we compared with.
B.2 Details on Calculating Surface Normal
@@
We define eigenvector with smallest eigenvalue as the Surface Normal $\mathbf{n}_{p}$. The defination is consistent with [Guédon and Lepetit, 2024]. However, $\mathbf{n}_{p}$ is unoriented, meaning it can point either inside or outside. Therefore, we determine the surface normal direction by considering the object center (in practice, we use proposal center). To elaborate, the orientation of surface normal will align with the orientation pointing from the object center to the Gaussian center. An alternative strategy is to follow [Ran et al., 2022] to determine the orientation. Specifically, we ensure that the first element of the normal vector is non-negative. Then we determine the orientation using an instance-level random inversion with $p=50\%$. Ablation study in Table 2 shows that the result of strategy of [Ran et al., 2022] is lower than the aforementioned strategy.
B.3 More Details on Training
@@
B.4 Proof of Theorem 1
Theorem 1 is a special case of Gauss’ Flux Theorem by setting $\mathbf{T}$ as any constant vector field:
$\oiint\limits_{S}\mathbf{T}\cdot\mathbf{n}\mathrm{d}S=\iiint\limits_{V}\nabla\cdot\mathbf{T}\mathrm{d}V\stackrel{{\scriptstyle\nabla\cdot\mathbf{T}=0}}{{=}}0$ (20)
Appendix C Additional Ablation Experiments and Analysis
C.1 Hyper-parameter Setup
@@
C.2 Impacts of $F_{V}$ on the Training Process
@@
C.3 Analysis on Underlying Gaussian Representation
Table 9: Quantitative results on the Gaussian representation.
Methods 3D-FRONT ScanNet
AP25 AP50 AP25 AP50
Gaussian-Det (3DGS+SuGaR [Kerbl et al., 2023, Guédon and Lepetit, 2024]) 96.7 77.7 71.7 24.5
Gaussian-Det (3DGS [Kerbl et al., 2023]) 94.8 74.3 68.9 22.0
Gaussian-Det (2DGS [Huang et al., 2024]) 95.3 79.9 70.1 24.6
To investigate how the underlying Gaussian representation, we have experimented on the ablation study by detecting from 3G Gaussians that were reconstructed via the original 3DGS [Kerbl et al., 2023] algorithm.
The quantitative results are displayed in Table B. While the detection accuracies drop with the input of lower quality 3D Gaussians, they still outnumber all the compared methods in Table 9 which consume higher-quality 3D Gaussians that were reconstructed via 3DGS+SuGaR. We additionally show experimental results using 2DGS for better surface modeling quality and improvements can be observed.
Appendix D Additional Visulization Results
D.1 Additional Qualitative Results on Ablation Studies
In Figure 11, we have accordingly illustrated the qualitative comparisons of the ablation studies on different components (see Figure 11 in the Appendix of the revised manuscript). It can be seen that either discarding the holistic surface coalescence or using the deterministic partial surface representation leads to sub-optimal visual results, such as inaccurate sizes and locations.
Figure 11: Additional qualitative results on ablation studies.
D.2 Additional Qualitative Results under Challenging Scenarios
@@
In Figure 13, we have included more qualitative results on occlusions, where the compared methods predict inaccurate sizes or locations.
Figure 13: Additional qualitative results on challenging cases and comparisons with other methods.
Another challenging scenario is detecting from low-quality 3D Gaussians that are reconstructed from images with noised poses. From Figure 14, we can see that Gaussian-Det is still capable of detecting the target objects, especially the table which is visually blurred after rendering.
Figure 14: Qualitative results on low-quality Gaussians reconstructed from images with noised poses.
D.3 Additional Failure Cases
@@
Appendix E Additional Quantitative Results
E.1 Detailed Comparison with NeRF-RPN and NeRF-MAE
@@
E.2 Model Efficiency on 3D Instance Segmentation
Apart from the quantitative results in Table 6, we have additional measured model efficiency in Table 11. We can see that leveraging the surface closure can substantially speed up the training period and significantly save the GPU memory footprint.
Table 11: Comparison of model efficiency on the task of 3D instance segmentation.
Methods Training Time GPU Memory
Gaussian-Grouping [Ye et al., 2024] 1.33h 35.8GB
Gaussian-Grouping + Ours 0.65h 12.8GB